# mFabric: An Efficient and Scalable Fabric for Mixture-of-Experts Training

### v2

@@ 5.3. Collective Communication Allocation
With the generated topology, the next step is to allocate network traffic from different parallelisms to mFabric and generate routing schedules. In the following, we illustrate how mFabric’s collective communications manager routes network traffic from various parallelisms in the data path.
TP and PP. TP traffic is limited to intra-host high bandwidth domain (e.g., the NVSwitch). PP traffic occurs across different PP stages and relies on the high-fanout EPS fabric in the mFabric architecture. As a result, there are no special configurations for these two types of parallelism.
DP. DP traffic typically spans the entire training cluster. Therefore, mFabric routes it through the EPS fabric. To further improve the communication efficiency of DP transfers, we leverage the hierarchical all-reduce algorithm (Wan et al., 2020; Jiang et al., 2020) to reduce the outbound traffic volume from each server. First, the GPUs within each server perform an intra-host reduction to aggregate the parameters to a gateway DP GPU connected to the EPS NIC. Next, all servers engage in a global ring all-reduce among the gateway DP GPUs to synchronize the model parameters. Finally, each server broadcasts the synchronized parameters from the gateway DP GPU to all other GPUs. The first and third stages of communication use the high-speed NVSwitch, while the second stage relies on the relatively lower-bandwidth EPS fabric. If multiple EPS NICs are available in the fabric, mFabric utilizes a multi-ring all-reduce method to fully exploit the bandwidth and reduce communication time.
Topology-aware EP. EP traffic is expected to use the regionally reconfigurable high-bandwidth domain. After reconfiguration, mFabric essentially forms a direct-connect topology for EP transfers. Based on that, mFabric uses the following steps to route the EP traffic. We illustrate this process in Figure 8, which depicts a reconfigured topology among four servers with a total of 16 GPUs (EP degree equals 16).
@@
    Each GPU looks up the topology to identify its intra-server communication delegation GPU for all communication pairs. mFabric prioritizes directly connected optical circuits over EPS. For example, the delegation GPU from server 0 to server 2 is GPU 2, as they are connected with optical circuits. However, to perform communications between server 0 and server 1, they have to use GPUs that are connected in the EPS.
@@
    With the gateway information, each server performs an intra-host gather, gathering outbound data to the corresponding delegation GPUs via NVSwitch. Note that in §5.2, we balanced the number of NICs across each NUMA node to mitigate intra-host congestion when multiple links are provisioned between a server pair. mFabric aims to distribute the traffic load across delegation NICs as evenly as possible.
@@
Figure 8. Routing of all-to-all communications in mFabric. For presentation simplicity, each server only contains 4 GPUs, and the TX/RX links are merged into a single link. In practice, mFabric supports the standard setup of 8 GPUs or more.
As the dataflows in steps (3) and (4) do not interfere with each other, mFabric overlaps the communication in these two steps to reduce overall completion time.
@@ 5.4. Failure Handling
There are two categories of failures in distributed MoE training: network (NIC/link) failures and GPU failures. mFabric is tolerant to both transient and permanent failures during distributed MoE training.
@@
Network fault resilience. In practice, each server in mFabric connects to both the global EPS fabric and a regional OCS domain using multiple NICs. Specifically, the EPS-side typically uses at least two NICs per server (see §4.2), providing redundancy for packet-switched communications. Based on prior reports (Qian et al., 2024), the failure rate of a single NIC or link during a training job is approximately 0.057%, making dual EPS NICs sufficient to reduce the probability of simultaneous failure to below 0.00003%. If both EPS NICs on a server fail—a rare but possible event—mFabric reroutes traffic through the OCS domain. Specifically, traffic destined for the failed NICs is first routed optically to a healthy peer, and then relayed through that peer’s functional EPS interface. Similarly, the EPS can serve as a fallback path if OCS links or ports fail. This dual-path design ensures resilient connectivity, at the cost of some additional intra-cluster forwarding overhead.
@@
GPU fault tolerance. mFabric also supports failure recovery when one or more GPUs become unavailable. We consider two realistic failure levels:
@@
  • Single-GPU failure. If a GPU fails during training, mFabric remaps the workload to a designated backup GPU. This is aligned with the design of high-availability systems such as NVL72, which reserve spare GPUs per group for fault tolerance. In mFabric, the backup GPU may be connected via either EPS or OCS: 1) If reachable via EPS, training resumes without additional routing. 2) If reachable only via OCS, mFabric forwards traffic through a peer GPU with optical connectivity to the backup GPU after topology reconfiguration, maintaining functional interconnect through minor topology adjustments.
@@
  • Full-node failure. A complete server failure (e.g., all 8 GPUs) requires a replacement node from the global backup pool. These backup nodes connect via EPS uplinks, ensuring network connectivity without reliance on regional OCS. Upon checkpoint restoration, mFabric resumes training with minimal disruption.
@@
Runtime reconfiguration. To maintain topology validity in the presence of failures, mFabric’s decentralized topology controllers detect communication failures and regenerate the OCS topology accordingly. This involves excluding failed nodes from the candidate set and recomputing optical mappings (see Algorithm 1 in §5.2). Because mFabric relies on regional control, such reconfigurations are localized and incur minimal global disruption. We evaluate the impact of various failure scenarios in §7.5.
6. mFabric Prototype
@@
To evaluate mFabric, we build a fully functional prototype using commodity hardware [^10] Due to Nvidia’s warranty restrictions, we cannot reconfigure the topology of the DGX SuperPod used in the measurement study (§3). Hence, we use commodity servers equipped with Nvidia GPUs for the testbed. capable of training state-of-the-art MoE models.
@@
Figure 9. mFabric testbed using commodity hardware.
Software stack. We implement mFabric’s software stack using approximately 6K lines of code in C++, including the topology generator, the OCS controller, and the custom collective communication runtime supporting in-training topology reconfigurations. For DP and PP communication over the static EPS fabric, we use NCCL (ncc, [n. d.]) to provide high-speed intra-host and inter-host all-reduce/point-to-point communications. For EP’s all-to-all communications that involve both EPS and OCS, our custom collective communication runtime leverages RDMA for high-speed data transfer using the raw ibverbs library. We port the mFabric runtime to Python to integrate with Megatron-LM (Shoeybi et al., 2020) for training real-world MoE models. Specifically, we have implemented communication primitives similar to those in torch.dist and expose them as mfabric.all_to_all and mfabric.all_reduce.
@@
Training state-of-the-art MoE models. We use the prototype to train three state-of-the-art MoE models, and compare its performance with a baseline configuration where all the four NICs are connected to an Ethernet switch (the ideal switch baseline). Figure 10 shows that mFabric achieves comparable performance to the 4 $\times$ 100G EPS baseline. mFabric utilizes one NIC in the EPS fabric and configures the remaining three NICs in an optical circuit fabric (a total of 12 optical ports and 4 electronic ports). In contrast, the EPS baseline uses the four 100 Gbps ConnectX-6 NICs in a non-blocking EPS fabric with 16 electronic ports. mFabric’s performance stems from its ability to efficiently provision high-bandwidth optical circuits for communication-intensive pairs in sparsely non-uniform all-to-all traffic, without compromising the transfer speed of DP and TP traffic. It is important to note that mFabric does not alter the parallelization strategies used in MoE training, but only optimizes data transfer through its architectural design and efficient circuit-switching algorithm. As a result, mFabric does not affect the training accuracy of MoE models.
7. Large-Scale Simulations
This section evaluates the performance of mFabric through large-scale simulation. We also present design space explorations on several factors, such as network scalability, EPS link options, and reconfiguration delays in Appendix §D.
@@ 7.1. Setup
Simulated GPU interconnect fabrics. We compare the performance of mFabric with the following interconnects:
  • mFabric (this work). In mFabric, each server connects two NICs to the EPS fabric using a fat-tree topology and connects the remaining six NICs to the OCS fabric by default. Following the architecture of the regionally reconfigurable high-bandwidth domain in mFabric, the optical circuit switch only needs to connect the GPUs within a single EP group, which is a maximum of 64 GPUs in our configuration. This can be easily supported by commodity OCS technologies (Table 1). mFabric blocks the network for 25 ms during the reconfiguration of the OCS for the first all-to-all communication in the forward pass and hides the reconfiguration time during computation for subsequent all-to-all communications, as discussed in §5.1.
@@
  • OverSub. Fat-tree. We compare mFabric with a Fat-tree interconnect with the 3:1 over-subscription ratio.
@@
7.2. Networking Cost Analysis
We present the cost analysis of mFabric in Figure 11, using a common production setup where each server contains 8 GPUs, following the same methodology as (Wang et al., 2023). The networking cost is analyzed with link bandwidths from 100 Gbps to 800 Gbps across different cluster sizes. It is important to note that we only account for the number of actually used switch ports in calculating the cost, as the cluster may not fit perfectly within a fat-tree/rail-optimized topology with a reasonable $K$. More details on the cost of each networking component can be found in Appendix D.2.
First, compared to the non-blocking Fat-tree and Rail-optimized topologies, mFabric reduces networking costs by an average of 2.0$\times$, as it organizes its high-bandwidth domain using OCS interconnects, which is significantly cheaper than EPS fabrics at high link bandwidth. Specifically, as shown in Figure 11(c), mFabric’s OCS fabric incurs 2.3$\times$ lower cost on average than fat-tree topology at 400 Gbps.
Second, we acknowledge that mFabric incurs slightly higher expenses than TopoOpt at the cluster size of 128 servers (1024 GPUs). This is because: 1) mFabric requires EPS fabric to maintain global network-wide connectivity, and 2) mFabric’s high-bandwidth domains assume millisecond-level reconfigurable OCS to adapt to runtime MoE traffic, which is more expensive than TopoOpt’s slowly reconfigurable patch panel. However, TopoOpt requires a multi-tier patch panel fabric to form a network of more than 1K GPUs. Achieving this requires extensive patch panel ports and expensive long-reach transceivers to compensate for the insertion loss of optical signals across multiple switching layers. As a result, it remains unclear whether TopoOpt is able to interconnect such large clusters and maintain its cost-efficiency.
7.3. Performance: Training Speed Ups
This section compares the end-to-end training iteration time of mFabric against other interconnects across four MoE models on the cluster with 128 servers and 1024 GPUs.
Figure 12(a) compares the training iteration time of various interconnects for the Mixtral 8$\times$22B model. Due to its efficient bandwidth allocation, we observe that mFabric achieves performance very close to the ideal Fat-tree and Rail-optimized topologies. In particular, with a TP degree of 8, mFabric provides direct optical circuits for almost all high-traffic server pairs during all-to-all communication (24 optical circuits for 8 EP participants). Compared to TopoOpt, mFabric reduces the training iteration time by 1.5$\times$ on average, as TopoOpt’s static topology cannot adapt to real-time traffic variations. In addition, we observe that mFabric outperforms the over-subscribed fat-tree by up to 1.6$\times$. Figure 12(b) shows a similar trend for the Mixtral 8$\times$7B model, with mFabric reducing the iteration time by 1.4$\times$ on average compared to TopoOpt. Both Mixtral models show diminishing returns from increased link bandwidth, as they are computation-bound at a micro-batch size of 8 (Figure 3). At higher bandwidths, the communication overhead shrinks, narrowing the performance gap between mFabric and others. The results for larger batch sizes of these Mixtral models are in the Appendix D.4.
Figure 12(c) and Figure 12(d) compare mFabric with other interconnects on Qwen-MoE (32 experts with 32-way EP) and DeepSeek-R1 (256 experts with 64-way EP), both of which use larger numbers of experts and higher EP degrees than Mixtral models. We observe that mFabric achieves performance comparable to Rail-optimized and Fat-tree topologies, and outperforms TopoOpt by 1.5$\times$ on average in Qwen-MoE and 1.3$\times$ in DeepSeek-R1. Compared to Fat-tree, mFabric exhibits slightly larger performance gaps at low bandwidths as the number of experts increases (e.g., 32 in Qwen-MoE vs. 8 in Mixtral 8$\times$22B). This is because more bandwidth-intensive GPU pairs require dedicated optical circuits, which slightly exceeds mFabric’s default optical fanout under the 8-GPU setup (6 circuits for 32 EP participants). Nevertheless, owing to the sparsity of EP traffic, mFabric narrows the performance gap as link bandwidth increases. We also observe that, unlike Qwen-MoE, mFabric shows smaller performance gains on DeepSeek-R1 as bandwidth increases. This is because, although DeepSeek-R1 uses a higher EP degree, its larger model size results in a lower communication-to-computation ratio, thus reducing the benefit of additional bandwidth.

### v3

@@ 5.3. Collective Communication Allocation
With the generated topology, the next step is to allocate network traffic from different parallelisms to MixNet and generate routing schedules. In the following, we illustrate how MixNet’s collective communications manager routes network traffic from various parallelisms in the data path.
TP and PP. TP traffic is limited to intra-host high bandwidth domain (e.g., the NVSwitch). PP traffic occurs across different PP stages and relies on the high-fanout EPS fabric in the MixNet architecture. As a result, there are no special configurations for these two types of parallelism.
DP. DP traffic typically spans the entire training cluster. Therefore, MixNet routes it through the EPS fabric. To further improve the communication efficiency of DP transfers, we leverage the hierarchical all-reduce algorithm (Wan et al., 2020; Jiang et al., 2020) to reduce the outbound traffic volume from each server. First, the GPUs within each server perform an intra-host reduction to aggregate the parameters to a gateway DP GPU connected to the EPS NIC. Next, all servers engage in a global ring all-reduce among the gateway DP GPUs to synchronize the model parameters. Finally, each server broadcasts the synchronized parameters from the gateway DP GPU to all other GPUs. The first and third stages of communication use the high-speed NVSwitch, while the second stage relies on the relatively lower-bandwidth EPS fabric. If multiple EPS NICs are available in the fabric, MixNet utilizes a multi-ring all-reduce method to fully exploit the bandwidth and reduce communication time.
Figure 8. Routing of all-to-all communications in MixNet. For presentation simplicity, each server only contains 4 GPUs, and the TX/RX links are merged into a single link. In practice, MixNet supports the standard setup of 8 GPUs or more.
Topology-aware EP. EP traffic is expected to use the regionally reconfigurable high-bandwidth domain. After reconfiguration, MixNet essentially forms a direct-connect topology for EP transfers. Based on that, MixNet uses the following steps to route the EP traffic. We illustrate this process in Figure 8, which depicts a reconfigured topology among four servers with a total of 16 GPUs (EP degree equals 16).
@@
    Each GPU looks up the topology to identify its intra-server communication delegation GPU for all communication pairs. MixNet prioritizes directly connected optical circuits over EPS. For example, the delegation GPU from server 0 to server 2 is GPU 2, as they are connected with optical circuits. However, to perform communications between server 0 and server 1, they have to use GPUs that are connected in the EPS.
@@
    With the gateway information, each server performs an intra-host gather, gathering outbound data to the corresponding delegation GPUs via NVSwitch. Note that in §5.2, we balanced the number of NICs across each NUMA node to mitigate intra-host congestion when multiple links are provisioned between a server pair. MixNet aims to distribute the traffic load across delegation NICs as evenly as possible.
@@
As the dataflows in steps (3) and (4) do not interfere with each other, MixNet overlaps the communication in these two steps to reduce overall completion time.
@@ 5.4. Failure Handling
There are two categories of failures in distributed MoE training: network (NIC/link) failures and GPU failures. MixNet is tolerant to both transient and permanent failures during distributed MoE training.
@@
Network fault resilience. In practice, each server in MixNet connects to both the global EPS fabric and a regional OCS domain using multiple NICs. Specifically, the EPS-side typically uses at least two NICs per server (see §4.2), providing redundancy for packet-switched communications. Based on prior reports (Qian et al., 2024), the failure rate of a single NIC or link during a training job is approximately 0.057%, making dual EPS NICs sufficient to reduce the probability of simultaneous failure to below 0.00003%. If both EPS NICs on a server fail—a rare but possible event—MixNet reroutes traffic through the OCS domain. Specifically, traffic destined for the failed NICs is first routed optically to a healthy peer, and then relayed through that peer’s functional EPS interface. Similarly, the EPS can serve as a fallback path if OCS links or ports fail. This dual-path design ensures resilient connectivity, at the cost of some additional intra-cluster forwarding overhead.
@@
GPU fault tolerance. MixNet also supports failure recovery when one or more GPUs become unavailable. We consider two realistic failure levels:
@@
  • Single-GPU failure. If a GPU fails during training, MixNet remaps the workload to a designated backup GPU. This is aligned with the design of high-availability systems such as NVL72, which reserve spare GPUs per group for fault tolerance. In MixNet, the backup GPU may be connected via either EPS or OCS: 1) If reachable via EPS, training resumes without additional routing. 2) If reachable only via OCS, MixNet forwards traffic through a peer GPU with optical connectivity to the backup GPU after topology reconfiguration, maintaining functional interconnect through minor topology adjustments.
@@
  • Full-node failure. A complete server failure (e.g., all 8 GPUs) requires a replacement node from the global backup pool. These backup nodes connect via EPS uplinks, ensuring network connectivity without reliance on regional OCS. Upon checkpoint restoration, MixNet resumes training with minimal disruption.
@@
Runtime reconfiguration. To maintain topology validity in the presence of failures, MixNet’s decentralized topology controllers detect communication failures and regenerate the OCS topology accordingly. This involves excluding failed nodes from the candidate set and recomputing optical mappings (see Algorithm 1 in §5.2). Because MixNet relies on regional control, such reconfigurations are localized and incur minimal global disruption. We evaluate the impact of various failure scenarios in §7.5.
6. MixNet Prototype
@@
To evaluate MixNet, we build a fully functional prototype using commodity hardware [^8] Due to Nvidia’s warranty restrictions, we cannot reconfigure the topology of the DGX SuperPod used in the measurement study (§3). Hence, we use commodity servers equipped with Nvidia GPUs for the testbed. capable of training state-of-the-art MoE models.
@@
Figure 9. MixNet testbed using commodity hardware.
Software stack. We implement MixNet’s software stack using approximately 6K lines of code in C++, including the topology generator, the OCS controller, and the custom collective communication runtime supporting in-training topology reconfigurations. For DP and PP communication over the static EPS fabric, we use NCCL (ncc, [n. d.]) to provide high-speed intra-host and inter-host all-reduce/point-to-point communications. For EP’s all-to-all communications that involve both EPS and OCS, our custom collective communication runtime leverages RDMA for high-speed data transfer using the raw ibverbs library based on FuseLink (Ren et al., 2025). We port the MixNet runtime to Python to integrate with Megatron-LM (Shoeybi et al., 2020) for training real-world MoE models. Specifically, we have implemented communication primitives similar to those in torch.dist and expose them as mixnet.all_to_all and mixnet.all_reduce.
Training state-of-the-art MoE models. We use the prototype to train three state-of-the-art MoE models, and compare its performance with a baseline configuration where all the four NICs are connected to an Ethernet switch (the ideal switch baseline). Figure 10 shows that MixNet achieves comparable performance to the 4 $\times$ 100G EPS baseline. MixNet utilizes one NIC in the EPS fabric and configures the remaining three NICs in an optical circuit fabric (a total of 12 optical ports and 4 electronic ports). In contrast, the EPS baseline uses the four 100 Gbps ConnectX-6 NICs in a non-blocking EPS fabric with 16 electronic ports. MixNet’s performance stems from its ability to efficiently provision high-bandwidth optical circuits for communication-intensive pairs in sparsely non-uniform all-to-all traffic, without compromising the transfer speed of DP and TP traffic. It is important to note that MixNet does not alter the parallelization strategies used in MoE training, but only optimizes data transfer through its architectural design and efficient circuit-switching algorithm. As a result, MixNet does not affect the training accuracy of MoE models.
7. Large-Scale Simulations
This section evaluates the performance of MixNet through large-scale simulation. We also present design space explorations on several factors, such as network scalability, EPS link options, and reconfiguration delays in Appendix §D.
@@ 7.1. Setup
Simulated GPU interconnect fabrics. We compare the performance of MixNet with the following interconnects:
  • MixNet (this work). In MixNet, each server connects two NICs to the EPS fabric using a fat-tree topology and connects the remaining six NICs to the OCS fabric by default. Following the architecture of the regionally reconfigurable high-bandwidth domain in MixNet, the optical circuit switch only needs to connect the GPUs within a single EP group, which is a maximum of 64 GPUs in our configuration. This can be easily supported by commodity OCS technologies (Table 2). MixNet blocks the network for 25 ms during the reconfiguration of the OCS for the first all-to-all communication in the forward pass and hides the reconfiguration time during computation for subsequent all-to-all communications, as discussed in §5.1.
@@
  • OverSub. Fat-tree. We compare MixNet with a Fat-tree interconnect with the 3:1 over-subscription ratio.
@@
7.2. Networking Cost Analysis
We present the cost analysis of MixNet in Figure 11, using a common production setup where each server contains 8 GPUs, following the same methodology as (Wang et al., 2023). The networking cost is analyzed with link bandwidths from 100 Gbps to 800 Gbps across different cluster sizes. It is important to note that we only account for the number of actually used switch ports in calculating the cost, as the cluster may not fit perfectly within a fat-tree/rail-optimized topology with a reasonable $K$. More details on the cost of each networking component can be found in Appendix D.2.
First, compared to the non-blocking Fat-tree and Rail-optimized topologies, MixNet reduces networking costs by an average of 2.0$\times$, as it organizes its high-bandwidth domain using OCS interconnects, which is significantly cheaper than EPS fabrics at high link bandwidth. Specifically, as shown in Figure 11(c), MixNet’s OCS fabric incurs 2.3$\times$ lower cost on average than fat-tree topology at 400 Gbps.
Second, we acknowledge that MixNet incurs slightly higher expenses than TopoOpt at the cluster size of 128 servers (1024 GPUs). This is because: 1) MixNet requires EPS fabric to maintain global network-wide connectivity, and 2) MixNet’s high-bandwidth domains assume millisecond-level reconfigurable OCS to adapt to runtime MoE traffic, which is more expensive than TopoOpt’s slowly reconfigurable patch panel. However, TopoOpt requires a multi-tier patch panel fabric to form a network of more than 1K GPUs. Achieving this requires extensive patch panel ports and expensive long-reach transceivers to compensate for the insertion loss of optical signals across multiple switching layers. As a result, it remains unclear whether TopoOpt is able to interconnect such large clusters and maintain its cost-efficiency.
7.3. Performance: Training Speed Ups
This section compares the end-to-end training iteration time of MixNet against other interconnects across four MoE models on the cluster with 128 servers and 1024 GPUs.
Figure 12(a) compares the training iteration time of various interconnects for the Mixtral 8$\times$22B model. Due to its efficient bandwidth allocation, we observe that MixNet achieves performance very close to the ideal Fat-tree and Rail-optimized topologies. In particular, with a TP degree of 8, MixNet provides direct optical circuits for almost all high-traffic server pairs during all-to-all communication (24 optical circuits for 8 EP participants). Compared to TopoOpt, MixNet reduces the training iteration time by 1.5$\times$ on average, as TopoOpt’s static topology cannot adapt to real-time traffic variations. In addition, we observe that MixNet outperforms the over-subscribed fat-tree by up to 1.6$\times$. Figure 12(b) shows a similar trend for the Mixtral 8$\times$7B model, with MixNet reducing the iteration time by 1.4$\times$ on average compared to TopoOpt. Both Mixtral models show diminishing returns from increased link bandwidth, as they are computation-bound at a micro-batch size of 8 (Figure 3). At higher bandwidths, the communication overhead shrinks, narrowing the performance gap between MixNet and others. The results for larger batch sizes of these Mixtral models are in the Appendix D.4.
Figure 12(c) and Figure 12(d) compare MixNet with other interconnects on Qwen-MoE (32 experts with 32-way EP) and DeepSeek-R1 (256 experts with 64-way EP), both of which use larger numbers of experts and higher EP degrees than Mixtral models. We observe that MixNet achieves performance comparable to Rail-optimized and Fat-tree topologies, and outperforms TopoOpt by 1.5$\times$ on average in Qwen-MoE and 1.3$\times$ in DeepSeek-R1. Compared to Fat-tree, MixNet exhibits slightly larger performance gaps at low bandwidths as the number of experts increases (e.g., 32 in Qwen-MoE vs. 8 in Mixtral 8$\times$22B). This is because more bandwidth-intensive GPU pairs require dedicated optical circuits, which slightly exceeds MixNet’s default optical fanout under the 8-GPU setup (6 circuits for 32 EP participants). Nevertheless, owing to the sparsity of EP traffic, MixNet narrows the performance gap as link bandwidth increases. We also observe that, unlike Qwen-MoE, MixNet shows smaller performance gains on DeepSeek-R1 as bandwidth increases. This is because, although DeepSeek-R1 uses a higher EP degree, its larger model size results in a lower communication-to-computation ratio, thus reducing the benefit of additional bandwidth.